solve the ratio of x+2 and 4 + 8 = x+4
Answer by dabanfield(803)

The ratio is (x+2)/(4+8) so we have:
(x+2)/12 = x+4
Multiply both sides by 12:
x+2 = 12x+48
Simplify:
11x = -46
Divide both sides by 11:
x=-46/11